OpenStudy (help!!!!):

8. A scale on a map states that every ¼ of an inch represents 20 miles. If two cities are 3 ½ inches apart, how many miles are actually between the two cities? A. 14 miles B. 20 miles C. 125 miles D. 230 miles E. 280 miles

OpenStudy (anonymous):

one inch would be 80 miles 3 inches would be 240 miles plus another 40 miles left over for a 1/2 in, so 240+40=

OpenStudy (jennychan12):

proportions. \[\frac{0.25 }{ 20 } = \frac{ 3.5 }{ x }\] solve for x.

OpenStudy (jennychan12):

0.25 x = 70 x = 280

OpenStudy (jennychan12):

E